OpenStudy (anonymous):

what is the area of the region bounded by the curves: y = x^{3} - x & y= 7x-x^{3}

OpenStudy (anonymous):

\[curves: y = x^{3} - x and y = 7x - x^{3}\]

OpenStudy (anonymous):

"I meant for there to be spaces between "and"

OpenStudy (anonymous):

Relatively easy - I just want to make sure I got the right answer: 16

OpenStudy (anonymous):

\[2x^3 = 7x\] \[\implies 2x^2 = 7\] \[\implies x= \sqrt{7/2}\] So the area between the curves is \[\int_0^{\sqrt{7/2}} [7-x^3 - x^3]dx = \int_0^{\sqrt{7/2}}[7-2x^3]dx\] \[=7x - \frac{2x^4}{4} |_0^{\sqrt{7/2}}\]

OpenStudy (anonymous):

I don't get 16.

OpenStudy (anonymous):

Oh whoops. I missed an x there.

OpenStudy (anonymous):

Should be \[=\frac{7x^2}{2} - \frac{2x^4}{4} |_0^{\sqrt{7/2}}\]

OpenStudy (anonymous):

But I still don't get 16